# Multi-Object Hallucination in Vision-Language Models

### v1

@@ 4.1 LVLM Baselines
  • LVLMs with base LLMs at different scales: LLaVA v1.6 (7B/13B/34B) (Liu et al., 2023b, 2024) and Yi-VL (6B/34B) (Young et al., 2024).
  • LVLMs with conversational/grounded instruction tuning: QwenVL-Base/Chat (7B) (Bai et al., 2023) and CogVLM-Base/Chat/Grounding v1.1 (19B) (Wang et al., 2023c).
  • Mechanistically grounded LVLMs: GLaMM (7B) (Rasheed et al., 2024) and GroundHOG (7B) (Zhang et al., 2024b).
  • Other LVLMs: IDEFICS-instruct (9B) (Laurençon et al., 2023b), MiniCPM-V v2.5 (8B) (Hu et al., 2024; Yu et al., 2024b), GPT-4V (OpenAI, 2023), and GPT-4O (OpenAI, 2024).
For mechanistically grounded LVLMs that take visual prompts through specially designed mechanisms, such as pointer tokens in GroundHOG (Zhang et al., 2024b), we additionally experiment with their default format and report whichever yields higher performance. For other LVLMs, we overlay the visual prompts on the images using a red bounding box with a width of 2 and visual text specifying the object index, presented with a white italic font on a black background with an alpha value of 0.75 for contrast and visibility.
@@ 4.2 Main Results and Findings
Our immediate observation is that LVLMs suffer from more hallucinations when tasked with focusing on and recognizing multiple objects compared to a single object. Across most of the models and test splits, we find that the average accuracy of single-object queries (i.e., probing object classes one at a time) significantly outperforms that of all three types of multi-object queries. The first exceptions are GPT-4O, MiniCPM-V, and CogVLM-2, the latter two leverage LLaMA-3 (Meta, 2024). Another exception to this is when teacher-forcing is applied to homogeneous test splits, which demonstrates an unreasonably high accuracy. We discuss them later in this section.
@@ 4.3 What May Help and What May Not?
We find that using base LLMs with more parameters reduces single-object hallucinations, but may not have the same effect on multi-object hallucinations. We observe a consistent increase in performance with larger LLaVA models in the seen set and in single-object queries, but not in the unseen set with multi-object queries. One possible explanation for this finding is that LLMs with more parameters are better at memorizing seen images, as the performance gap between seen and unseen images is also more significant in larger models. We also notice that the performance gap between single-object probing and multi-object probing does not apply to MiniCPM-V and CogVLM-2, which adopt a LLaMA-3 (8B) (Meta, 2024) base LLM pre-trained with 15T tokens, as they fail to follow the instruction sometimes. Compared to LLaVA models developed upon LLaMA-2 (7/13B) (Touvron et al., 2023) and Yi (34B) (Young et al., 2024) with 2T and 3T pre-training tokens, these models underperform in quantitative measures due to instruction following error but exhibit greater robustness when multiple visual prompts are presented.
@@
While it’s surprising that conversational tuning reduces multi-object hallucinations, we observe that models without conversational tuning struggle to follow instructions and are prone to shortcuts, such as repeating the list of all object class candidates in order or consistently repeating the first candidate. This might also explain why grounded tuning in CogVLM-G is of little help in reducing multi-object hallucinations thus far. These models typically lack conversational fine-tuning, and there is currently no available grounded dialogue data at scale. While mechanistically grounded LVLMs show strong performances in single-object probing, there remain a gap in multi-object probing with student forcing. This could be attributed to a significant portion of the grounded instruction tuning dataset consisting mainly of short captions or questions featuring one single or few objects.
@@ 5.1 Potential Hallucinatory Factors
  • Object Token Position: Zhou et al. (2024) has shown that more hallucinations occur in the latter part of captions. In this work, the object indices directly correspond to the object token positions.
@@
  • Object Centrality: Previous research has identified a center bias in datasets and models, indicating that objects are disproportionately located at the center of images in detection models (Szabó and Horváth, 2022; Taesiri et al., 2023). We define object centrality as the distance $d$ between the object’s bounding box center and the image center, normalized by the diagonal distance $D$ from the center to the corner.
@@
  • Object Salience: Previous research has shown that smaller objects are harder to detect and ground to (Hoiem et al., 2012; Ma et al., 2023). We define object salience as the ratio of the number of pixels occupied by the object’s instance segmentation mask to the total number of pixels in the image.
  • Training Salience: Previous research has shown that spurious co-occurring patterns in the training data can lead to object hallucinations (Li et al., 2023a; Zhou et al., 2024). We use the log frequency of classes in MSCOCO as a proxy for training salience following previous work, and hypothesize that LVLMs tend to hallucinate more on less frequent objects in the training set.
Model Behaviors.
  • Object Token Entropy: Zhou et al. (2024) have shown that object hallucinations are more likely when the decoded object tokens have a higher log perplexity. In our work, we define object token entropy as the entropy of the logits of the first token in the generated word. Given $\mathbf{s}$ as the softmax logits of the generated word’s first token, we calculate the entropy using the following formula: $H(\mathbf{s})=-\sum_{i}\mathbf{s}_{i}\log(\mathbf{s}_{i}).$ Simply put, higher entropy indicates greater uncertainty in the model’s prediction for the first token, which can lead to more frequent object hallucinations.
  • Visual Modality Contribution: We hypothesize that LVLMs pay less attention to the visual modality during object hallucinations. Motivated by the modality importance score (Cao et al., 2020), we define Visual Modality Contribution (VMC) as the proportion of attention allocated to visual tokens compared to textual tokens. To quantify this, we analyze the attention weights of the last generated token across all heads and layers. The VMC is computed as follows: $\textrm{VMC}=\sum_{i\in V\alpha_{ij}}/\big{(}\sum_{i\in V}\alpha_{ij}+\sum_{k\in T}\alpha_{kj}\big{)}$, where $\alpha_{ij}$ represents the attention weight assigned to visual token $i$ at head $j$, and $\alpha_{kj}$ represents the attention weight assigned to textual token $k$ at head $j$. The sets $V$ and $T$ denote the visual and textual tokens, respectively. By examining the VMC, we can determine how much attention is given to visual inputs in comparison to textual inputs. A lower VMC may indicate a higher likelihood of object hallucinations due to insufficient attention to visual cues.
5.2 When Do LVLMs Experience Multi-Object Hallucinations?
Data/Task-specific Factors.
Salience and Frequency.
@@ 5.2 When Do LVLMs Experience Multi-Object Hallucinations?
The intrinsic behaviors of the model provide significant insights into its tendencies to hallucinate. Similar to Zhou et al. (2024), we find that models are more prone to hallucination when they experience uncertainty or confusion, especially in scenarios involving multiple objects, as evidenced by higher token entropy. Furthermore, the contribution from the visual modality consistently registers below 20%, suggesting that current LVLMs may rely more heavily on linguistic contexts. There is a marginal increase in the likelihood of hallucination when models pay less attention to the visual context.
@@ Acknowledgement
This work was supported in part by NSF IIS-1949634, NSF SES-2128623, and the DARPA Machine Common Sense Program. Our experiments have also benefited from the Microsoft Accelerate Foundation Models Research (AFMR) grant program and the model access from the Amazon AGI team. The authors would like to thank Yichi Zhang for his valuable feedback.
Limitations
Ethical concerns and risks
@@ A.1 Reproducibility
Our data curation pipeline involves several essential steps designed to prepare and refine our dataset for evaluating multi-object hallucination. The pipeline begins by filtering images and candidate objects to query. We consider valid objects to be those belonging to the top 50 “thing” classes and exclude objects with a bounding box area less than 1% of the total image area. We discard images containing fewer than 5 valid objects, and allow an intersection-over-union between bounding boxes of no more than 0.1, which preserves data integrity while ensuring high image quality. We apply this pipeline to MSCOCO-Panoptic (Lin et al., 2014; Caesar et al., 2018) and ADE20K (Zhou et al., 2017). [^2] Available at https://huggingface.co/datasets/sled-umich/ROPE
Dataset Total COCO ADE
Wild 1539 / 1172 732 / 547 807 / 625
Hom. 312 / 490 168 / 289 144 / 201
Het. 400 / 246 200 / 76 200 / 170
Adv. 168 / 334 54 / 170 114 / 164
Table 3: Data statistics (Train/Validation)
@@
Artifacts and licenses
@@
Table 4: License information for the scientific artifacts.
A.2 Additional Experiments and Results
We provide the per-object performance in the following tables.

### v2

@@ 4.1 LVLM Baselines
  • LVLMs with base LLMs at different scales: LLaVA v1.6 (7B/13B/34B) (liu2023visual; liu2023improved) and Yi-VL (6B/34B) (young2024yi).
  • LVLMs with conversational/grounded instruction tuning: QwenVL-Base/Chat (7B) (bai2023qwen) and CogVLM-Base/Chat/Grounding v1.1 (19B) (wang2023cogvlm).
  • Mechanistically grounded LVLMs: GLaMM (7B) (rasheed2023glamm) and GroundHOG (7B) (zhang2024groundhog).
  • LVLMs with RL-based finetuning: MiniCPM-V (Yu2023RLHFVTT)
  • Other LVLMs: IDEFICS-instruct (9B) (laurenccon2024obelics), MiniCPM-V v2.5 (8B) (hu2024minicpm; yu2024rlaif), GPT-4V (gpt4v), and GPT-4O (gpt4o).
For mechanistically grounded LVLMs that take visual prompts through specially designed mechanisms, such as pointer tokens in GroundHOG (zhang2024groundhog), we additionally experiment with their default format and report whichever yields higher performance. For other LVLMs, we overlay the visual prompts on the images using a red bounding box with a width of 2 and visual text specifying the object index, presented with a white italic font on a black background with an alpha value of 0.75 for contrast and visibility.
@@ 4.2 Main Results and Findings
Our immediate observation is that LVLMs suffer from more hallucinations when tasked with focusing on and recognizing multiple objects compared to a single object. Across most of the models and test splits, we find that the average accuracy of single-object queries (i.e., probing object classes one at a time) significantly outperforms that of all three types of multi-object queries. The first exceptions are GPT-4O, MiniCPM-V, and CogVLM-2, the latter two leverage LLaMA-3 (llama3). Another exception to this is when teacher-forcing is applied to homogeneous test splits, which demonstrates an unreasonably high accuracy. We discuss them later in this section.
@@ 4.3 What May Help and What May Not?
We find that using base LLMs with more parameters reduces single-object hallucinations, but may not have the same effect on multi-object hallucinations. We observe a consistent increase in performance with larger LLaVA models in the seen set and in single-object queries, but not in the unseen set with multi-object queries. One possible explanation for this finding is that LLMs with more parameters are better at memorizing seen images, as the performance gap between seen and unseen images is also more significant in larger models. We also notice that the performance gap between single-object probing and multi-object probing does not apply to MiniCPM-V and CogVLM-2, which adopt a LLaMA-3 (8B) (llama3) base LLM pre-trained with 15T tokens, as they fail to follow the instruction sometimes. Compared to LLaVA models developed upon LLaMA-2 (7/13B) (touvron2023llama) and Yi (34B) (young2024yi) with 2T and 3T pre-training tokens, these models underperform in quantitative measures due to instruction following error but exhibit greater robustness when multiple visual prompts are presented.
@@
While it’s surprising that conversational tuning reduces multi-object hallucinations, we observe that models without conversational tuning struggle to follow instructions and are prone to shortcuts, such as repeating the list of all object class candidates in order or consistently repeating the first candidate. This might also explain why grounded tuning in CogVLM-G is of little help in reducing multi-object hallucinations thus far. These models typically lack conversational fine-tuning, and there is currently no available grounded dialogue data at scale. While mechanistically grounded LVLMs show strong performances in single-object probing, there remain a gap in multi-object probing with student forcing. This could be attributed to a significant portion of the grounded instruction tuning dataset consisting mainly of short captions or questions featuring one single or few objects.RL-based finetuning approaches, such as MiniCPM-V, demonstrate promising results across diverse settings, surpassing single-object results in both the Wild and Homogenous settings. Upon inspection, we found that this model demonstrates strong visual in-context learning capability and improves correct recognition when objects of the same classes are probed together.
Decoding and inference time strategy
Decoding algorithms like OPERA introduce nuanced improvements in specific multi-object settings(Huang2023OPERAAH). In default multi-object tasks, OPERA shows marginal performance enhancements for LLaVA-1.5, but its effectiveness declines in tests with greater heterogeneity, to the point it can even lower performance. This suggests OPERA is beneficial in homogeneous contexts but requires further refinement in handling mixed object scenarios effectively.
@@ 5.1 Potential Hallucinatory Factors
  • Object Token Position: zhou2023analyzing has shown that more hallucinations occur in the latter part of captions. In this work, the object indices directly correspond to the object token positions.
@@
  • Object Centrality: Previous research has identified a center bias in datasets and models, indicating that objects are disproportionately located at the center of images in detection models (szabo2022mitigating; taesiri2023imagenet). We define object centrality as one minus the distance $d$ between the object’s bounding box center and the image center, normalized by the diagonal distance $D$ from the center to the corner.
Salience and Frequency.
  • Object Salience: Previous research has shown that smaller objects are harder to detect and ground to (hoiem2012diagnosing; ma2023world). We define object salience as the ratio of the number of pixels occupied by the object’s instance segmentation mask to the total number of pixels in the image.
  • Training Salience: Previous research has shown that spurious co-occurring patterns in the training data can lead to object hallucinations (li2023evaluating; zhou2023analyzing). We use the log frequency of classes in MSCOCO as a proxy for training salience following previous work, and hypothesize that LVLMs tend to hallucinate more on less frequent objects in the training set.
Model Behaviors.
  • Object Token Entropy: zhou2023analyzing have shown that object hallucinations are more likely when the decoded object tokens have a higher log perplexity. In our work, we define object token entropy as the entropy of the logits of the first token in the generated word. Given $\mathbf{s}$ as the softmax logits of the generated word’s first token, we calculate the entropy using the following formula: $H(\mathbf{s})=-\sum_{i}\mathbf{s}_{i}\log(\mathbf{s}_{i}).$ Simply put, higher entropy indicates greater uncertainty in the model’s prediction for the first token, which can lead to more frequent object hallucinations.
  • Visual Modality Contribution: We hypothesize that LVLMs pay less attention to the visual modality during object hallucinations. Motivated by the modality importance score (cao2020behind), we define Visual Modality Contribution (VMC) as the proportion of attention allocated to visual tokens compared to textual tokens. To quantify this, we analyze the attention weights of the last generated token across all heads and layers. The VMC is computed as follows: $\textrm{VMC}=\sum_{i\in V\alpha_{ij}}/\big{(}\sum_{i\in V}\alpha_{ij}+\sum_{k\in T}\alpha_{kj}\big{)}$, where $\alpha_{ij}$ represents the attention weight assigned to visual token $i$ at head $j$, and $\alpha_{kj}$ represents the attention weight assigned to textual token $k$ at head $j$. The sets $V$ and $T$ denote the visual and textual tokens, respectively. By examining the VMC, we can determine how much attention is given to visual inputs in comparison to textual inputs. A lower VMC may indicate a higher likelihood of object hallucinations due to insufficient attention to visual cues.
5.2 When Do LVLMs Experience Multi-Object Hallucinations?
Data/Task-specific Factors.
@@ 5.2 When Do LVLMs Experience Multi-Object Hallucinations?
The intrinsic behaviors of the model provide significant insights into its tendencies to hallucinate. Similar to zhou2023analyzing, we find that models are more prone to hallucination when they experience uncertainty or confusion, especially in scenarios involving multiple objects, as evidenced by higher token entropy. Furthermore, the contribution from the visual modality consistently registers below 20%, suggesting that current LVLMs may rely more heavily on linguistic contexts. There is a marginal increase in the likelihood of hallucination when models pay less attention to the visual context.
@@ Acknowledgement
This work was supported in part by NSF IIS-1949634, NSF SES-2128623, the DARPA Perceptual Task Guidance (PTG) Program, and the DARPA Machine Common Sense Program. Our experiments have also benefited from the Microsoft Accelerate Foundation Models Research (AFMR) program. We thank the Amazon AGI team for GroundHOG model access. The authors would like to thank Yichi Zhang and anonymous reviewers for their valuable feedback.
@@ A.1 Reproducibility
{wraptable}
!tr0.5 Dataset Total COCO ADE Wild 1539 / 1172 732 / 547 807 / 625 Hom. 312 / 490 168 / 289 144 / 201 Het. 400 / 246 200 / 76 200 / 170 Adv. 168 / 334 54 / 170 114 / 164
Our data curation pipeline involves several essential steps designed to prepare and refine our dataset for evaluating multi-object hallucination. The pipeline begins by filtering images and candidate objects to query. We consider valid objects to be those belonging to the top 50 “thing” classes and exclude objects with a bounding box area less than 1% of the total image area. We discard images containing fewer than 5 valid objects, and allow an intersection-over-union between bounding boxes of no more than 0.1, which preserves data integrity while ensuring high image quality. We apply this pipeline to MSCOCO-Panoptic (lin2014microsoft; caesar2018coco) and ADE20K (zhou2017scene). [^2] Available at https://huggingface.co/datasets/sled-umich/ROPE
@@
Figure 8: Single and multi-object hallucination under the default setting in nuScenes (caesar2020nuscenes).
@@
A.2 Additional Experiments and Results
Per-object Performance.
We provide the per-object performance in the following Table B.3 and Table B.3.
Autonomous Driving Case Study.
Figure 8 is a case study example from the nuScenes dataset (caesar2020nuscenes) for autonomous driving. It illustrates the single and multi-object case, where each object is identified independently. The multi-object case exhibits more hallucination errors compared to the single-object case. This finding underscores the importance of studying multi-object hallucination, especially in real-world scenarios like autonomous driving, where multiple objects need to be detected accurately at the same time.
Appendix B Limitations, Licenses, and Risks
B.1 Limitations
B.2 Artifacts and licenses
@@
Table 3: License information for the scientific artifacts used.
B.3 Ethical concerns and risks
@@
Table 2: An overview of object hallucination benchmarks. For design considerations, we summarize the number of tested images, and if multiple classes and object class distribution (at training and test time) are considered. The image sources include those seen or unseen during instruction tuning. To refer to an object, textual descriptions and visual cues can be adopted. For evaluation, neural models, humans and automatic pipelines are used.